A solubility of a salt is 80g in 100g of water at 100 degree Celsius and 20g in 100g of water at 25 degree Celsius. What mass of salt would crystallize if 50g of the salt solution saturated at 100 degree Celsius were cooled at 25 degree Celsius



1
Expert's answer
2022-01-19T04:02:02-0500

Mass of solution at 100°C "=100+80"

"=180g"


Mass of solution at 25°C"=100+20"

"=120g"


Mass of salt that crystalize when 180g of solution is cooled from 100°C to 30°C

"=80-20"

"=60g"


When 50g of solution is cooled from 100°C to 30°C


Mass of salt that crystalizes

"=\frac{50}{180}×60"


"=16.67g"